OpenStudy (anonymous):

A tourist calculated that if he walks to the railroad station with a speed of 4mph, he'll miss the train by half an hour, but if he walks with a speed of 5 mph, he'll reach the station 6 min before the departure of the train. What distance does the tourist have to cover?

OpenStudy (mertsj):

Remember that rate times time = distance. They have given us two different expressions for the distance tp the train station? The first one is 4(x+.5) Can you figure out the second one?

OpenStudy (anonymous):

5(x-0.1) @Mertsj

OpenStudy (mertsj):

Yes. And those two distances must be equal

OpenStudy (anonymous):

5(x-0.1) = 4(x+0.5) 5x-0.5 = 4x + 2 x = 2.5

OpenStudy (mertsj):

You are very smart!!

OpenStudy (anonymous):

@Mertsj Thanks Mr.Genius!

OpenStudy (mertsj):

yw

OpenStudy (dan815):

:)

OpenStudy (dan815):

ask another one!!